# Quantum Computing for Query Containmentof Conjunctive Queries

### v1

@@ A.2. Simulated annealing & quantum annealing
After describing QAOA in the previous subsection, we now present the foundations of two further optimization methods that have been used to solve the QUBOs considered in this work. These methods are referred to as simulated annealing and quantum annealing (see also Subsection 2.2).
@@
where $\sigma_{i}^{x}$ is the Pauli-$x$ operator for qubit $q_{i}$. Applying $H_{I}$ yields a superposition of the different quantum states in $Q$ such that in the ground state of $H_{I}$ (energy minimizing state) every quantum state in $Q$ is equally likely to be observed. Combining $H_{I}$ and $H_{p,\textit{an}}$, we obtain a Hamiltonian depending on a variable $s$ parameterized by $t$:
@@
Adapting $A$ and $B$, allows us to change between different minima of the energy landscape by tunneling through energy barriers that separate them. If the energy barriers are thin, the probability of tunneling increases. In contrast to simulated annealing, the probability governing the change between local minima is not affected by the difference between the minima. Exploiting the tunneling phenomenon, we can move faster and further across the energy landscape compared to the simulated annealing method, since within a single iteration of the simulated annealing method, one has to “climb” over energy barriers according to the Boltzmann distribution instead of tunneling through the energy barriers. During the annealing process, the energy landscape is more and more adapted to the considered QUBO. If $A$ and $B$ are modified sufficiently slowly, the whole process will finish in an energy minimizing ground state with high probability due to the adiabatic theorem. Since $A$ and $B$ change gradually, this method can be assigned to the class of annealing based optimization methods. The whole method is summarized in Algorithm 2.
@@ Appendix C Details omitted in Section 4
We can find a homomorphism from an $i$-chain to a 2-cycle by “wrapping” the chain around the 2-cycle. Depending on where we start this (left or right node of the 2-cycle), we obtain two different possibilities. This provides two distinct homomorphisms that solve the containment problem for every $i\in\mathbb{N}_{\geqslant 1}$. More formally, the first homomorphism $h$ is given via $h(Y_{k})=Z_{0}$ if $k$ is even and $h(Y_{k})=Z_{1}$ if $k$ is odd (for $0\leqslant k\leqslant i$). The second homomorphism $h^{\prime}$ is defined the other way round, via $h^{\prime}(Y_{k})=Z_{1}$ if $k$ is even and $h^{\prime}(Y_{k})=Z_{0}$ if $k$ is odd (for $0\leqslant k\leqslant i$). Each of these two individual solutions can be specified by the mapping of the first vertex $Y_{0}$ of the $i$-chain to either $Z_{0}$ or $Z_{1}$ of the 2-cycle, because this uniquely determines the rest of the homomorphism. Mapping the other vertices follows accordingly. See Figure 10(a) for a visualization, where the two mappings are indicated in blue and yellow. Since we can find a homomorphism from the $i$-chain to the 2-cycle, the latter is contained in the former (but not vice versa). The ground truth is given by these two homomorphisms. Their bit-matrix representations have i+1i{+}1 rows $Y_{0},\ldots,Y_{i}$ and $2$ columns $Z_{0},Z_{1}$; any two consecutive rows are of the form 10 and 01; and the bit-matrix for $h$ starts with row 10, while the bit-matrix for $h^{\prime}$ starts with row 01. Each of these two bit-matrices is represented by the bitvector obtained by concatenating the rows of the matrix, i.e., $h$ is represented by the bitvector $10011001\cdots$, and $h^{\prime}$ is represented by the bitvector $01100110\cdots$.
@@ Appendix E Details omitted in Section 6
Since the algorithms regarded in this paper are heuristic, their performance depends on the individual problem structure. While they have a certain expected performance on average, problem families can be constructed to precisely fit or not fit the nature of the algorithm. Such differences can become more apparent when artificial examples are considered. The course of our algorithms is influenced by the so-called “solution landscape” or “energy landscape” that is explored during the run. Recall that we represent a query pair $(q_{1},q_{2})$, for which we want to decide whether or not $q_{1}\sqsubseteq q_{2}$ holds, by a polynomial $p(\overline{x})$. Every possible value of $p(\overline{x})$ for an assignment of the binary variables $\overline{x}$ with 0-1-values, can be regarded as an “energy level” or “value” of an “objective function”. Thereby, every possible solution candidate (i.e., assignment of $\overline{x}$ with 0-1-values; henceforth they will be called “bitvectors”) corresponds to a point on the energy landscape.
The overall goal is to reach the global minimum by an annealing-inspired process. As explained in Section 2.2 and Section A.2, the simulated annealing algorithm strives for states of lower or similar energy while sometimes (with exponentially decreasing probability) exploring neighboring states of higher energy than the current one (see the definition of $AcceptProb$ in Section A.2).
As we will see, the experimental results can be explained by differences in the distribution and connectedness of energy levels, so let us have a closer look at the corresponding energy landscapes. We first give an overview of the visible macro-effects and follow with a more detailed computation of the individual relevant sizes.
@@
Here, we focus on $(q_{1},q_{2})$ being either the family of query pairs 2-cycle to $i$-chain or the family of query pairs 2-chain to $i$-star, for $i\in\mathbb{N}_{\geqslant 1}$. For both families, the energy levels can be partitioned into a very large portion of positive-valued states representing invalid candidates (pos states), a number of zero-valued states representing incorrect mappings with respect to the tuples in $q_{2}$ (zero states), and an amount of negative-valued states representing at least partially correct mappings (neg states). The relative size of pos states approaches 100% as $i$ grows, and there is an increasing energy gap of 2i+12i{+}1 to the zero and neg states, which is defined by the chosen “penalty weight” $|\textbf{T}_{1}|{\cdot}|\textbf{T}_{2}|+1$. For the small portion of zero and neg states, it is worth investigating the development of their proportions towards each other. While both families start off with a relatively large zero plateau, its fraction declines towards 0 for 2-cycle to $i$-chain and approaches a 50/50 balance with the neg states for 2-chain to $i$-star. A summary for the development of all ratios is given in Table 3. In order to interpret these ratios with respect to the experimental results, we also need to investigate the connectedness of the states.
@@
For 2-cycle to $i$-chain, the negative energy levels are connected in a cascading/short-stepped way, such that connected neg states are at most two energy levels away from each other. This is due to the chain structure of the tuples in $q_{2}$. Every node of $q_{2}$ is part of at most two edges (and thus of at most two tuples), which means that changing one single bit can change at most two terms in $p_{\textup{(3)}}(\overline{x})$. Therefore, reaching an optimal solution from the zero level takes at least $\frac{i}{2}$ steps. Moreover, the energy landscape contains a number of local minima that are not optimal, which are indicated by crosses in the figure. They exist because independent parts of each of the two optimal solutions can contribute to a negative value within the same bitvector. Such a local but not global minimum could, for example, look like $LRLRLRRLRLLR$, where $L$ stands for $10$ (mapping a node from the chain to the left node $Z_{0}$) and $R$ for $01$ (mapping it to the right node $Z_{1}$). The example candidate consists of three parts ($LRLRLR;RLRL;LR$) that independently look like part of an optimal solution and only the connections between the parts are off. In general, x-candidates look like piece-wise optimal candidates: sub-chains of size at least 2 are mapped correctly in itself (alternating between left and right) but alternate between the two optimal solutions. There is no way to decrease the value of that state without first increasing it by “unmapping” one of the nodes. These local minima occur below energy levels of $\frac{i}{2}$, and their number increases with growing $i$, which plays a role in lowering the solution probabilities for that family.
@@
For 2-chain to $i$-star, the neg states can be partitioned into two separate parts, half of which contain bitvectors starting with ‘100’ (central node $Y_{0}$ mapped to $Z_{0}$, indicated in blue) and half of which start with ‘010’ (central node $Y_{0}$ mapped to $Z_{1}$, indicated in yellow). This is due to the central node of the star being part of every single tuple and thus of every negative term in the polynomial. Moreover, this means that there are no local minima except for the two global minima. Every neg state can reach the next lower energy level without increasing its energy by at most two steps. If one of the nodes in $\{Y_{1},\ldots,Y_{i}\}$ has not been mapped correctly, it is either not mapped at all or mapped incorrectly. In the latter case, it can be unmapped without increasing the energy level and then be correctly mapped in the next step. Regarding the first case, it can directly be mapped to reach the lower energy level. The absence of local-only minima is a key factor for the good performance observed in the experiments, because the algorithm cannot get stuck.
@@
It remains to clarify why we observe lower solution probabilities in the beginning. This can be explained as follows: The neg states are highly interconnected with the zero states, since every neg state has a direct zero neighbor where the first 1 is replaced by 0 (starting with ‘000’, so the central node is unmapped). This holds specifically for the two optimal states, which also have a zero neighbor. As explained before, there is a low chance to jump out of the optimum into that higher-energy zero neighbor. For simulated annealing, we can directly compute that probability, which decreases exponentially as $i$ grows and annealing time passes by (see Equation (7)). The counterprobability of staying in the optimum therefore increases logarithmically. Figure 12 illustrates that the probability to “not escape” for growing $i$ at $25\%$ of time and compares it to the measured solution probability in the experiment. Our hypothesis is that this point in time is close to the critical point when the escape probability is still high enough to have an impact but afterwards quickly vanishes. This is backed up by the high similarity of the two scatter plots. In combination with the relatively large zero plateau for small $i$, jumping out of the optimum can cause that the algorithm gets lost afterwards, since the transition probability between states of the same energy level is always 1 and thus the algorithm can move to other zero states after escaping the optimum. Putting it all together, this explains the lower solution probabilities observed for small $i$ and their logarithmic growth as $i$ increases. For exact computations of the probabilities depicted in Figure 12, see the end of the next paragraph.
@@
Figure 12. Solution probability for 2-chain-to-$i$-star vs. the probability to not escape the optimal solution to its zero neighbor at 25% of the time for growing $i$
@@
The relative sizes of the different state portions are computed in the following. A state has a positive value if $p_{\textup{fct}}(\overline{x})>0$, i.e., if the bit-matrix contains a row with at least two ones. For 2-cycle to $i$-chain, each row has length 2, so in three out of four possible 2-bit row-vectors, the row as at most one 1. The portion of pos states is given by the counterprobability of all rows having at most one 1:
@@
In the case of 2-cycle to $i$-chain, the state is negative if at least two consecutive nodes from the $i$-chain are mapped to opposite nodes from the 2-cycle. Abstractly, we can think of having an alphabet $\Sigma:=\{0,L,R\}$ and the language $\mathbf{L}_{i}:=\Sigma^{i+1}\cap(*(LR|RL)*)$ containing all words of length i+1i{+}1 that contain the sub-word $LR$ or $RL$, which represents the portion of neg states. The size of that language can be computed (see https://oeis.org/A193519; accessed 2025/12/02) and the resulting fraction is given by
$p_{i}^{(\text{chain})}(\text{neg}|\neg\text{pos})=\frac{1}{2^{i+2}}(2{\cdot}3^{i+1}-(1+\sqrt{2})^{i+2}-(1-\sqrt{2})^{i+2})$
which approaches 1. Therefore, the opposite portion of zero states approaches 0.
In the case of 2-chain to $i$-star, a state has a negative value if the central node $Y_{0}$ of the $i$-star is mapped to either $Z_{0}$ or $Z_{1}$ of the 2-chain, and if at least one of the other nodes is mapped to $Z_{1}$ or $Z_{2}$, respectively. Given a correct mapping of the central node, each other node has a probability of $\frac{3}{4}$ to not follow suit. So the fraction of zero states, given that each row has at most one 1, can be computed by
@@
$p_{i}^{(\text{star})}(neg|\neg\text{pos})\ \ =\ \ \textstyle\frac{1}{2}\left(1-\left(\frac{3}{4}\right)^{i}\right).$
@@
After understanding how to compute the numbers in Table 3, we explain the computations required to understand Figure 12. We assume that an optimal state has been reached and want to compute the probability to escape into the zero plateau when the algorithm encounters the zero neighbor. Following Algorithm 1, this means that $x$ is the optimal state and $x_{\text{new}}$ is the zero neighbor. For energy levels, it follows c=ic=-i and $c_{\text{new}}=0$, so the probability of transition follows AcceptProb(T,i)(T,i) with the given cooling schedule $T$. In the code implementation, the schedule’s min and max are given by the values $\beta_{\text{start}}=\frac{1}{T_{\text{high}}k_{b}}:=0.5$ and $\beta_{\text{end}}=\frac{1}{T_{\text{low}}k_{b}}:=10$ and the algorithm follows a geometric schedule divided into 1000 steps, which means that for $k\in\mathbb{N}$ with $1\leqslant k\leqslant 1000$, we can compute
$\alpha=\sqrt[1000]{\frac{\beta_{\text{end}}}{\beta_{\text{start}}}}\quad\text{and}\quad\beta_{k}=\beta_{\text{start}}\cdot\alpha^{k}.$
@@
$\text{AcceptProb}_{.25}(T,i)=e^{-\beta_{250}\cdot i}=e^{-0.5\cdot\sqrt[4]{20}\cdot i}$
Taking the counterprobability $1-e^{-\beta_{250}\cdot i}$ we have finally arrived at the probability of staying in the optimal solution after 25% of the annealing time depending on $i$, which is shown in Figure 12.

### v2

@@ A.2. Simulated annealing & quantum annealing
After describing QAOA in the previous subsection, we now present the foundations of two further optimization methods that have been used to solve the QUBOs considered in this work. These methods are referred to as simulated annealing and quantum annealing (see also Section 2.2).
@@
where $\sigma_{i}^{x}$ is the Pauli-$x$ operator for qubit $\mathbf{q}_{i}$. Applying $H_{I}$ yields a superposition of the different quantum states in $R$ such that in the ground state of $H_{I}$ (energy minimizing state) every quantum state in $R$ is equally likely to be observed. Combining $H_{I}$ and $H_{p,\textit{an}}$, we obtain a Hamiltonian depending on a variable $s$ parameterized by $t$:
@@
Adapting $A$ and $B$ allows us to change between different minima of the energy landscape by tunneling through energy barriers that separate them. If the energy barriers are thin, the probability of tunneling increases. In contrast to simulated annealing, the probability governing the change between local minima is not affected by the difference between the minima. Exploiting the tunneling phenomenon, we can move faster and further across the energy landscape compared to the simulated annealing method, since within a single iteration of the simulated annealing method, one has to “climb” over energy barriers according to the Boltzmann distribution instead of tunneling through the energy barriers. During the annealing process, the energy landscape is more and more adapted to the QUBO considered. If $A$ and $B$ are modified sufficiently slowly, the whole process will finish in an energy minimizing ground state with high probability due to the adiabatic theorem. Since $A$ and $B$ change gradually, this method can be assigned to the class of annealing based optimization methods. The whole method is summarized in Algorithm 2.
@@ Appendix C Details omitted in Section 4
We can find a homomorphism from an $i$-chain to a 2-cycle by “wrapping” the chain around the 2-cycle. Depending on where we start this (left or right node of the 2-cycle), we obtain two different possibilities. This provides two distinct homomorphisms that solve the containment problem for every $i\in\mathbb{N}_{\geqslant 1}$. More formally, the first homomorphism $h$ is given via $h(Y_{k})=Z_{0}$ if $k$ is even and $h(Y_{k})=Z_{1}$ if $k$ is odd (for $0\leqslant k\leqslant i$). The second homomorphism $h^{\prime}$ is defined the other way round, via $h^{\prime}(Y_{k})=Z_{1}$ if $k$ is even and $h^{\prime}(Y_{k})=Z_{0}$ if $k$ is odd (for $0\leqslant k\leqslant i$). Each of these two individual solutions can be specified by the mapping of the first vertex $Y_{0}$ of the $i$-chain to either $Z_{0}$ or $Z_{1}$ of the 2-cycle, because this uniquely determines the rest of the homomorphism; mapping the other vertices follows accordingly. See Figure 10(a) for a visualization, where the two mappings are indicated in blue and yellow. Since we can find a homomorphism from the $i$-chain to the 2-cycle, the latter is contained in the former (but not vice versa). The ground truth is given by these two homomorphisms. Their bit-matrix representations have i+1i{+}1 rows $Y_{0},\ldots,Y_{i}$ and $2$ columns $Z_{0},Z_{1}$; for any two consecutive rows, one is of the form 10 and the other is of the form 01; and the bit-matrix for $h$ starts with row 10, while the bit-matrix for $h^{\prime}$ starts with row 01. Each of these two bit-matrices is represented by the bitvector obtained by concatenating the rows of the matrix, i.e., $h$ is represented by the bitvector $10011001\cdots$, and $h^{\prime}$ is represented by the bitvector $01100110\cdots$.
@@ Appendix E Details omitted in Section 6
The actual performance of the algorithms considered in this paper depends on the individual problem structure. While they have a certain expected performance on average, problem families can be constructed to precisely fit or not fit the nature of the algorithm. Such differences can become more apparent when artificial examples are considered. The course of our algorithms is influenced by the energy landscape that is explored during the run. Recall that we represent a query pair $(q_{1},q_{2})$, for which we want to decide whether $q_{1}\sqsubseteq q_{2}$ holds, by a polynomial $p(\overline{x})$. Every possible value of $p(\overline{x})$ for an assignment of the binary variables $\overline{x}$ with 0-1-values, can be regarded as an “energy level” or “value” of the objective function $p(\overline{x})$. Thus, every possible solution candidate (i.e., assignment of $\overline{x}$ with 0-1-values; henceforth they will be called “bitvectors”, “states” or “candidates”) corresponds to a point on the energy landscape.
The overall goal is to reach the global minimum by an annealing-inspired process. As explained in Section 2.2 and Section A.2, the simulated annealing algorithm strives for states of lower or similar energy while sometimes (with exponentially decreasing probability) exploring neighboring states of higher energy than the current one (see the definition of AcceptProb in Equation (8) of Section A.2).
As we will see, the experimental results can be explained by differences in the distribution and connectedness of energy levels, so let us take a closer look at the corresponding energy landscapes. We first give an overview of the visible macro-effects and follow with a more detailed computation of the individual relevant sizes.
@@
Here, we focus on $(q_{1},q_{2})$ being either the family of query pairs 2-cycle to $i$-chain or the family of query pairs 2-chain to $i$-star, for $i\in\mathbb{N}_{\geqslant 1}$. Our goal is to gain a better understanding of the observed scaling behavior depicted in Figure 5. For both families, the energy levels can be partitioned into a very large portion of positive-valued states representing invalid candidates, where at least one row of the represented matrix has more than one bit set to 1 (pos states), a number of zero-valued states representing incorrect mappings with respect to the tuples in $q_{2}$ (zero states), and an amount of negative-valued states representing at least partially correct mappings (neg states). The relative size of pos states approaches 100% as $i$ grows, and there is an increasing energy gap of 2i+12i{+}1 to the zero and neg states, which is defined by the chosen “penalty weight” $|\textbf{T}_{1}|{\cdot}|\textbf{T}_{2}|+1$. For the small portion of zero and neg states, it is worth investigating the development of their proportions towards each other. While both families start off with a relatively large zero plateau, its fraction declines towards 0 for 2-cycle to $i$-chain and approaches a 50/50 balance with the neg states for 2-chain to $i$-star. A summary of the development of all ratios is given in Table 3. In order to interpret these ratios with respect to the experimental results, we also need to investigate the connectedness of the states.
@@
For 2-cycle to $i$-chain, the negative energy levels are connected in a cascading/short-stepped way, such that connected neg states are at most two energy levels away from each other. This is due to the chain structure of the tuples in $q_{2}$. Every node of $q_{2}$ is part of at most two edges (and thus of at most two tuples), which means that changing one single bit can change at most two terms in $p_{\textup{(3)}}(\overline{x})$. Therefore, reaching an optimal solution from zero level takes at least $\frac{i}{2}$ steps. Moreover, the energy landscape contains a number of local minima that are not optimal, which are indicated by crosses in the figure. They exist because independent parts of each of the two optimal solutions can contribute to a negative value within the same bitvector. Such a local but not global minimum could, for example, look like $LRLRLRRLRLLR$, where $L$ stands for $10$ (mapping a node from the chain to the left node $Z_{0}$) and $R$ for $01$ (mapping it to the right node $Z_{1}$). The example candidate consists of three parts ($LRLRLR;RLRL;LR$) that independently look like part of an optimal solution, and only the connections between the parts are off. In general, those locally minimal (but not globally minimal) candidates look like piece-wise optimal candidates: sub-chains of size at least 2 are mapped correctly in itself (alternating between left and right) but alternate between the two optimal solutions. There is no way to decrease the value of that state without first increasing it by “unmapping” one of the nodes. These local minima occur below the energy levels of $\frac{i}{2}$, and their number increases with growing $i$, which plays a role in lowering the solution probabilities for that family.
@@
For 2-chain-to-$i$-star, the neg states can be partitioned into two separate parts, half of which contain bitvectors starting with ‘100’ (central node $Y_{0}$ mapped to $Z_{0}$, indicated in blue) and half of which start with ‘010’ (central node $Y_{0}$ mapped to $Z_{1}$, indicated in yellow). This is because the central node of the star is part of every single tuple and thus of every negative term in the polynomial. Moreover, this means that there are no local minima except for the two global minima. Every neg state can reach the next lower energy level without increasing its energy by at most two steps. If one of the nodes in $\{Y_{1},\ldots,Y_{i}\}$ has not been mapped correctly, it is either not mapped at all or incorrectly mapped. In the latter case, it can be unmapped without increasing the energy level and then correctly mapped in the next step. Regarding the first case, it can be directly mapped to reach the lower energy level. The absence of local-only minima is a key factor for the good performance observed in the experiments because the algorithm cannot get stuck.
@@
It remains to clarify why we observe lower solution probabilities for SA in the beginning, i.e., for small values of $i$ (cf., Figure 5(b)). This can be explained as follows: The neg states are highly interconnected with the zero states, since every neg state has a direct zero neighbor where the first 1 is replaced by 0 (starting with ‘000’, so the central node is unmapped). This holds specifically for the two optimal states, which also have a zero neighbor. As explained before, there is a low chance to jump out of the optimum into that higher-energy zero neighbor. For simulated annealing, we can directly compute that probability, which decreases exponentially as $i$ grows and annealing time passes by (see Equation (8)). The counterprobability of staying in the optimum is of the form $1-e^{-\gamma i}$ for a constant $\gamma>0$ (details on this are provided at the end of this section), which quickly approaches 1 for growing $i$. Figure 12 illustrates the probability to stay in the optimum for growing $i$ at $25\%$ of time and compares it to the measured solution probability of SA in the experiment that we showed in Figure 5(b). Our hypothesis is that after 25% of the time we are close to the critical point when the “escape probability” (i.e., the probability to escape from an optimal solution) is still high enough to have an impact, but afterward quickly vanishes. This is backed up by the correlation between the two scatter plots shown in Figure 12. In combination with the relatively large zero plateau for small $i$, jumping out of the optimum can cause the algorithm to get lost afterward, since the transition probability between states of the same energy level is always 1 (because then we have $c-c_{\text{new}}=0$ and $e^{0}=1$, cf. Equation (8)) and thus the algorithm can move to other zero states after escaping the optimum. Putting it all together, this explains the lower solution probabilities observed for small $i$ and their growth as $i$ increases. For exact computations of the probabilities depicted in Figure 12, see the end of the next paragraph.
@@
Figure 12. Measured solution probability (using SA simulator) vs. the theoretical probability for 2-chain-to-$i$-star to stay in optimal solution at 25% of the annealing time
@@
The relative sizes of the different state portions are computed as follows. A state $\overline{x}$ has a positive value if $p_{\textup{fct}}(\overline{x})>0$, i.e., if the bit-matrix contains a row with at least two ones. For 2-cycle to $i$-chain, each row has length 2, so in three out of four possible 2-bit row-vectors, the row as at most one 1. The portion of pos states is given by the counterprobability of all rows having at most one 1:
@@
In the case of 2-cycle to $i$-chain, the state is negative if at least two consecutive nodes from the $i$-chain are mapped to opposite nodes from the 2-cycle. Abstractly, we can think of having an alphabet $\Sigma:=\{N,L,R\}$ and the language $\mathbf{L}_{i}:=\Sigma^{i+1}\cap(*(LR|RL)*)$ containing all words of length i+1i{+}1 that contain the sub-word $LR$ or $RL$, which represents the portion of neg states. The size of that language can be computed (see https://oeis.org/A193519; accessed 2025/12/02) and is given by
$|\mathbf{L}_{i}|\ \ =\ \ {\textstyle\frac{1}{2}}\cdot\Big(2{\cdot}3^{i+1}-(1{+}\sqrt{2})^{i+2}-(1{-}\sqrt{2})^{i+2}\Big).$
The resulting fraction of all words of length i+1i{+}1 is then given by
$p_{i}^{(\text{chain})}(\text{neg}|\neg\text{pos})\ =\ \frac{1}{2{\cdot}3^{i+1}}\cdot\Big(2{\cdot}3^{i+1}-(1{+}\sqrt{2})^{i+2}-(1{-}\sqrt{2})^{i+2}\Big)$
which can also be written as $p_{i}^{(\text{chain})}(\text{neg}|\neg\text{pos})\ =$
[…]
and approaches 1 as $i$ increases. Therefore, the opposite portion of zero states approaches 0.
In the case of 2-chain to $i$-star, a state has a negative value if either the central node $Y_{0}$ of the $i$-star is mapped to $Z_{0}$ and at least one of the other nodes is mapped to $Z_{1}$, or if $Y_{0}$ is mapped to $Z_{1}$ and at least one of the other nodes is mapped to $Z_{2}$. Given a correct mapping of the central node, each other node has a probability of $\frac{3}{4}$ to not follow suit. So the fraction of zero states, given that each row has at most one 1, can be computed by
@@
$p_{i}^{(\text{star})}(\text{neg}|\neg\text{pos})\ \ =\ \ \textstyle\frac{1}{2}\left(1-\left(\frac{3}{4}\right)^{i}\right).$
@@
This explains how we computed the numbers in Table 3. Next, we explain the computations required to understand Figure 12.
We assume that an optimal state has been reached and want to compute the probability to escape into the zero plateau when the algorithm encounters a zero neighbor. Following Algorithm 1, this means that $\overline{x}$ is an optimal state and $\overline{x}_{\text{new}}$ is a zero neighbor. For their energy levels, we have c=ic=-i and $c_{\text{new}}=0$, so the probability of transition can be computed by $\textit{AcceptProb}(T,i)$ with the given cooling schedule $T$. In the code implementation, the schedule’s min and max are given by the values $\beta_{\text{start}}=\frac{1}{T_{\text{high}}k_{b}}:=0.5$ and $\beta_{\text{end}}=\frac{1}{T_{\text{low}}k_{b}}:=10$ and the algorithm follows a geometric schedule divided into 1000 steps, which means that for $\ell\in\mathbb{N}$ with $1\leqslant\ell\leqslant 1000$, we can compute
$\alpha=\sqrt[1000]{\frac{\beta_{\text{end}}}{\beta_{\text{start}}}}\quad\text{and}\quad\beta_{\ell}=\beta_{\text{start}}\cdot\alpha^{\ell}.$
@@
$\textit{AcceptProb}(T_{25\%},i)\ \ =\ \ e^{-\beta_{250}\cdot i}\ \ =\ \ e^{-0.5\cdot\sqrt[4]{20}\cdot i},$
where $T_{25\%}$ indicates using $\beta$ at 25% of the time, which is (due to using 1000 steps) given by $\beta_{250}$. Taking the counterprobability $1-e^{-\beta_{250}\cdot i}$ we have finally arrived at the probability of staying in the optimal solution after 25% of the annealing time depending on $i$, which is shown in Figure 12.